Suppose that sound is emitted uniformly in all directions by a public address system. The intensity at a location 26 m away from the sound source is 3.40 10-4 W/m2. What is the intensity at a spot that is 84 m away?
Thanks!!!
Answer #1

I2/I1 = (r1/r2)^2

I2/(3.4*10^-4) = (26/84)^2

I2 = 3.26*10^-5 W/m^2